the length of a rectangle is 5 inches more than the width. if the perimeter of the rectangle is 98 inches, what is the lenght and width?
Answer by checkley75(3666)

L=W+5
P=2L+2W
98=2(W+5)+2W
98=2W+10+2W
4W=98-10
4W=88
W=88/4
W=22 ANSWER.
L=22+5
L=27 ANSWER.
PROOF
98=2*22+2*27
98=44+54
98=98